OpenStudy (rebeccapink6398):

Twenty-five percent of what number is 7.5?

OpenStudy (anonymous):

Let the required number be x Therefore 25% of x = 7.5 25/100 * x= 7.5 25x=7.5*100 25x= 750 x= 750/25 x= 30 Hence the required number is 30. @RebeccaPink6398

OpenStudy (rebeccapink6398):

TIPE!! IT WAS CORRECT :D

OpenStudy (rebeccapink6398):

YIPPEEE* LOL that's what I mean